OpenStudy (anonymous):

solve the following matrix equation for a, b, c, d. [a - b b + c ] = [8 1] [3d + c 2a - 4d] [7 6] the ans : a = 5 b = -3 c = 4 d = 1

OpenStudy (anonymous):

please help

myininaya (myininaya):

so we have a-b=8 (eq1) b+c=1 (eq2) 3d+c=7 (eq3) 2a-4d=6 (eq4) --------------------------------- from the first equation we can write a=8+b We can plug this into equation 4 2(8+b)-4d=6 => 16+2b-4d=6 => 2b-4d=6-16 => 2b-4d=-10 => b-2d=-5 => b=-5+2d We can plug this into equation 2 (-5+2d)+c=1=>2d+c=6 => c=6-2d Finally plug this into equation 3 and solve for d

OpenStudy (anonymous):

is that the answer?

myininaya (myininaya):

no

myininaya (myininaya):

that will give you what d is

myininaya (myininaya):

but we haven't found all the other values a,b, and c

OpenStudy (anonymous):

i see

OpenStudy (anonymous):

how does a - b = 8 turn into a = 8 + b?

OpenStudy (anonymous):

ok nvm dumb question

myininaya (myininaya):

so I said c=6-2d and i asked you to plug this into equation 3 and equation 3 was 3d+c=7 (eq3) which gives 3d+(6-2d)=7 6+d=7 => d=1

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

I think i can get the rest

OpenStudy (anonymous):

THANK YOU VERY MUCH

myininaya (myininaya):

ok cool!